1) An open-top box is to be constructed from a 4 foot by 6 foot rectangular cardboard by cutting out equal squares at each corner and the folding up the flaps. Let x denote the length of each side of the square to be cut out.
a) Find the function V that represents the volume of the box in terms of x.
Answer:

Answer by checkley71(8403)

V=(6-2X)(4-2X)X
V=X(24-8X-12X+4X^2)
V=X(4X^2-20X+24)
V=4X(X^2-5X+6)
V=4X(X-2)(X-3)